OpenStudy (anonymous):

the volume v of a container is 84 ft^3.the width the length,and the height are x,x +1,x-4 respectively.what are the container's dimensions? please help me asap!!!!!!!!!!thanksss

OpenStudy (anonymous):

v = LWH plug in what you know and solve for x use the value of x to find LWH based on the respective info provided

OpenStudy (anonymous):

\[84=x(x+1)(x-4)\]

OpenStudy (anonymous):

foil the equation then find x

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

by trial set x=6 then x+1=7 x-4=2 6*7*4=84